# Robust Latent Representation Tuning for Image-text Classification

### v1

@@ Abstract
Large models have demonstrated exceptional generalization capabilities in computer vision and natural language processing. Recent efforts have focused on enhancing these models with multimodal processing abilities. However, addressing the challenges posed by scenarios where one modality is absent remains a significant hurdle. In response to this issue, we propose a robust latent representation tuning method for large models. Specifically, our approach introduces a modality latent translation module to maximize the correlation between modalities. Following this, a newly designed fusion module is employed to facilitate information interaction between the modalities. In this framework, not only are common semantics refined during training, but the method also yields robust representations in the absence of one modality. Importantly, our method maintains the frozen state of the image and text foundation models to preserve their abilities acquired through large-scale pretraining. We conduct experiments on several public datasets, and the results underscore the effectiveness of our proposed method.
Index Terms—  Image-text classification, large models, robust learning, representation learning.
@@ 1 Introduction
In recent times, large models have garnered substantial attention due to their remarkable generalization capabilities across numerous downstream tasks. Given that most large models are pretrained on unimodal datasets (e.g., LLaMA [1], OPT [2]), researchers have sought to augment these models with multimodal processing capabilities. Notably, approaches like LISA [3] have proposed extracting multimodal features using various large models, employing these features for tasks such as image segmentation. PixelLM [4] has introduced a tuning framework wherein visual embeddings are prefixed to textual tokens, jointly processed by large language models (LLM). Despite the numerous endeavors to imbue large models with the capacity to process multimodal signals (e.g., images and texts), there has been limited attention to robust representation learning, and performance in modality-absence scenarios remains relatively unexplored. Real-world applications frequently encounter scenarios where certain modalities are absent, rendering current methods challenging to apply.
To address this challenge, we introduce a novel strategy for robust multimodal representation tuning in this paper. Our approach leverages two pretrained large models dedicated to image and text processing. At each corresponding layer of the paired image-text models, we incorporate a Modality Latent Translation (MoLT) module. Within this module, image and text embeddings are projected onto a shared latent space, aiming to bring the embeddings closer together. This shared space acts as a bridge connecting the image and text domains. Subsequently, a cross-attention mechanism is employed after feature extraction to capture the relationship between the robust representation and the associated modality embeddings for making predictions.
@@
2 Related Prior Works
@@ 2.2 Multimodal Large Model Tuning
Fig. 1: The overview of our proposed method. The image and text are first processed by separate encoders for robust representation learning. After that we fuse the modality features and robust embedding for the final predictions.
@@ 3.1 Modality Latent Translation
When the image and text are processed by respective large models, a modality latent translation module is introduced to learn the robust representation. For each pair of image-text foundation model layers $l$, the corresponding representations are $I_{l}\in R^{N_{i}\times d_{i}}$ and $T_{l}\in R^{N_{t}\times d_{t}}$, where $N$ and $d$ are respective token numbers and dimensions. Then two linear projections are employed to map the embeddings into the same space:
@@
where $F=R_{11}^{-0.5}R_{12}R_{22}^{-0.5}$. Throughout the tuning process, the representations become more robust as the canonical correlation increases. Consequently, we can effectively translate representations from one modality to another, thereby empowering the model with the ability to infer in scenarios where one modality is missing.
@@ 3.2 Fusion and Training Target
After the frozen foundation model, the robust representations ($H_{i,l}^{{}^{\prime}}$ and $H_{t,l}^{{}^{\prime}}$) and extracted modality embeddings($E_{i}$, and $E_{t}$) are sent to our new designed fusion. The detailed structure is shown in Figure 1. We first introduce a learnable vector $M\in R^{L_{s}}$ to average-pool the robust representations from each selected image-text layers:
$H_{l}$ $=\frac{1}{2}(H_{i,l}^{{}^{\prime}}+H_{t,l}^{{}^{\prime}}),$ (5)
$H_{r}$ $=Avg(M\cdot[H_{1},H_{2},...,H_{l},...,H_{L_{s}}]),$
where $H_{l}$ is the joint robust representation of layer $l$ and $L_{s}$ is the number of selected layers for robust representation learning. Then we perform the information exchange between $H_{r}$, $E_{i}$, and $E_{t}$:
@@
where $Q_{r}$ is projected by $H_{r}$, $K_{i}$/$V_{i}$ are projected from $E_{i}$, and $K_{t}$/$V_{t}$ are projected from $E_{t}$. Finally, we utilize the mean of both for predictions:
$\hat{y}=Classifier(\frac{1}{2}(E_{r,i}+E_{r,t})).$ (7)
@@
where $L_{CE}$ is the cross entropy loss, $\alpha$ and $\beta$ are two hyper-parameters.
@@ 4.2 Experimental Settings
In our experiments, we employ the pretrained LLaMA as the text foundation model and the image encoder of CLIP-L/224 as visual foundation model. Inherited from the pretrained models, $d_{i}$ and $d_{t}$ are set to 4096 and 768, respectively. The dimension of common space $d_{c}$ is set to 1024. We infuse the MoLT module in the last 4 layers of image and text models, meaning that $L_{s}$ is 4. In the loss function, we set $\alpha$ to 0.1 and $\beta$ to 0.9. To reduce the memory consumption, we train our model with mixed-precision. The Adam optimizer is employed in our method and the learning rate is set to 0.0004. Our approach is implemented with PyTorch framework and the experiments are conducted on two NVIDIA RTX 3090Ti GPUs.
Table 1: The quantitative results of our method on three benchmark datasets. w/ LM indicates whether the large models are utilized in the approach.
@@
Ours $\checkmark$ 64.9/59.0 92.12 75.10
@@
Baseline $\checkmark$ 59.0 / 51.2 85.2 69.3
Ours(text-absence) $\checkmark$ 62.4 / 56.7 88.9 73.2
Ours(image-absence) $\checkmark$ 63.1 / 57.0 89.2 71.0
Table 2: The ablation study on SNLI-VE dataset. C.A. means the cross-attention in MoLT. Fusion indicates the fusion strategy in our method.
Ablation SNLI-VE
C.A. $L_{CCA}$ $M$ Fusion Acc(%)
$\checkmark$ $\checkmark$ $\checkmark$ 69.6
$\checkmark$ $\checkmark$ $\checkmark$ 70.5
$\checkmark$ $\checkmark$ $\checkmark$ 73.0
$\checkmark$ $\checkmark$ $\checkmark$ 71.9
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 73.45
Fig. 2: The visualization of some cases for our propose method and the baseline model.
4.3 Quantitative Results
The results of our method on the evaluated datasets are presented in Table 1. As evident from the results, we achieve state-of-the-art performance on each benchmark. Among the methods we compare with, HUSE [17] and VisualBert [18] do not utilize large models, while others (such as MaPLe [20], MMBT [19], and PMF [22]) are based on large models. Most large model-based methods aim to facilitate information exchange through fine-tuning. The performance gap observed among them underscores the potent generalization ability of large models.
4.4 Ablation Study
To further investigate the effectiveness of each component in our method, we conducted an ablation study using the SNLI-VE dataset. The results are presented in Table 2. When removing $L_{CCA}$, the performance drops dramatically, highlighting the crucial role of the training target. The results also demonstrate that the cross-attention module and the learnable vector $M$ have a positive influence on the final outcomes. When the fusion module is not utilized, meaning only the robust representations are employed for predictions, the performance also decreases. The performance gap observed between each ablative model and our final model underscores the effectiveness of our proposed method.
@@ 4.5 Robust Inference Analysis
As our focus lies in enhancing model robustness in modality-missing scenarios, we conducted corresponding experiments. Table 1 also presents the performance when inferring with only one modality. The results show that each modality has a varying impact on different tasks. For movie classification, the textual modality dominates, while the importance of the image modality increases for visual-entailment understanding. Nevertheless, the performance is consistently better than the baseline, which does not utilize the MoLT module and does not incorporate robust representation learning, thereby revealing the effectiveness of our proposed approach. Additionally, we visually inspect some cases in Figure 2. With the MoLT module and robust representation learning, the model can still predict results accurately, whereas the baseline model often fails.
@@ 5 Conclusion
In this paper, we propose a robust representation learning strategy tailored for large models. Our approach incorporates a modality latent translation module capable of translating one modality embedding to another. Additionally, we introduce a novel fusion schema for robust representation and modality embeddings. The experiments are conducted on three datasets, and the results clearly illustrate the effectiveness of our proposed method. In the future, we plan to conduct further research in robust representation learning to enhance our ability to handle modality-absence scenarios more effectively.

### v2

@@ Abstract
Large models have demonstrated exceptional generalization capabilities in computer vision and natural language processing. Recent efforts have focused on enhancing these models with multimodal processing abilities. However, addressing the challenges posed by scenarios where one modality is absent remains a significant hurdle. In response to this issue, we propose a robust latent representation tuning method for large models. Specifically, our approach introduces a modality latent translation module to maximize the correlation between modalities, resulting in a robust representation. Following this, a newly designed fusion module is employed to facilitate information interaction between the modalities. Within this framework, common semantics are refined during training, and robust performance is achieved even in the absence of one modality. Importantly, our method maintains the frozen state of the image and text foundation models to preserve their capabilities acquired through large-scale pretraining. We conduct experiments on several public datasets, and the results underscore the effectiveness of our proposed method.
Keywords:
Image-text classification large models robust learning representation learning
@@ 1 Introduction
In recent times, large models have garnered substantial attention due to their remarkable generalization capabilities across numerous downstream tasks. Given that most large models are pretrained on unimodal datasets (e.g., LLaMA [20], OPT [24]), researchers have sought to augment these models with multimodal processing capabilities. Notably, approaches like LISA [11] have proposed extracting multimodal features using various large models, employing these features for tasks such as image segmentation. PixelLM [18] has introduced a tuning framework wherein visual embeddings are prefixed to textual tokens, jointly processed by large language models (LLM). Despite the numerous endeavors to imbue large models with the capacity to process multimodal signals (e.g., images and texts), there has been limited attention to robust representation learning, and performance in modality-absence scenarios remains relatively unexplored. Real-world applications frequently encounter scenarios where certain modalities are absent or noised, rendering current methods challenging to apply.
To address this challenge, we introduce a novel strategy for robust multimodal representation tuning in this paper. Our approach leverages two pretrained large models dedicated to image and text processing. At each corresponding layer of the paired image-text models, we incorporate a Modality Latent Translation (MoLT) module. Within this module, image and text embeddings are projected onto a shared latent space for cross-modality interaction. This shared space acts as a bridge connecting the image and text domains. Subsequently, we employ factorized bilinear pooling (FBP)[23] to obtain the robust representation, which has been proven effective in multimodal learning[19]. After feature extraction, a cross-attention mechanism is employed to capture the relationship between the robust representation and the associated modality embeddings for making predictions.
@@
2 Related Works
In this section, we introduce some recent works related to our method, including large models and corresponding tuning method.
@@ 2.2 Multimodal Large Model Tuning
Figure 1: The overview of our proposed method. The image and text are first processed by separate encoders for robust representation learning. During this process, the robust representation $H_{r}$ is obtained by MolT module and FBP. After that we fuse the modality features and robust embedding for the final predictions.
@@ 3.1 Modality Latent Translation
When the image and text are processed by respective large models, a modality latent translation module(MolT) is introduced to learn the robust representation. For each pair of image-text foundation model layers $l$, the corresponding representations are $I_{l}\in R^{N_{i}\times d_{i}}$ and $T_{l}\in R^{N_{t}\times d_{t}}$, where $N$ and $d$ are respective token numbers and dimensions. Then two linear projections are employed to map the embeddings into the same space:
@@
where $F=R_{11}^{-0.5}R_{12}R_{22}^{-0.5}$. Throughout the tuning process, the representations become more robust as the canonical correlation increases. Consequently, we can effectively translate representations from one modality to another, thereby empowering the model with the ability to infer in scenarios where one modality is missing. Finally, we obtain the robust representation of layer $l$ by factorized bilinear pooling(FBP):
$H_{l}^{{}^{\prime}}$ $=SumPooling(H_{i,l}^{{}^{\prime}}\odot H_{t,l}^{{}^{\prime}},s),$ (5)
$H_{l}$ $=\frac{H_{l}^{{}^{\prime}}}{||H_{l}^{{}^{\prime}}||_{2}},$
where $\odot$ means the element-wise multiply and $s$ is the stride of sum pooling. The robust representation $H_{l}$ in each layer are utilized in the fusion stage for robust predictions.
@@ 3.2 Fusion and Training Target
After the frozen foundation model, the robust representations ($H_{l}$) and extracted modality embeddings($E_{i}$, and $E_{t}$) are sent to our new designed fusion. The detailed structure is shown in Figure 1. We first introduce a learnable vector $M\in R^{L_{s}}$ to average-pool the robust representations from each selected image-text layers:
$H_{r}$ $=Avg(M\cdot[H_{1},H_{2},...,H_{l},...,H_{L_{s}}]),$ (6)
where $L_{s}$ is the number of selected layers for robust representation learning. Then we perform the information exchange between $H_{r}$, $E_{i}$, and $E_{t}$:
@@
where $Q_{r}$ is projected by $H_{r}$, $K_{i}$/$V_{i}$ are projected from $E_{i}$, and $K_{t}$/$V_{t}$ are projected from $E_{t}$. To avoid the instability caused by the absence of a modality, we apply layer normalization before making the final predictions:
$\hat{y}=Classifier(\frac{1}{2}Norm(E_{r,i}+E_{r,t})).$ (8)
@@
where $L_{CE}$ is the cross entropy loss, $\alpha$ and $\beta$ are two hyper-parameters to balance the numerical scales of different losses.
@@ 4 Experiments and Analysis
In this section, we introduce our experimental settings, evaluated benchmarks and respective results.
@@ 4.2 Experimental Settings
In our experiments, we employ the pretrained LLaMA as the text foundation model and the image encoder of CLIP-L/224 as visual foundation model. Inherited from the pretrained models, $d_{i}$ and $d_{t}$ are set to 4096 and 768, respectively. The dimension of common space $d_{c}$ is set to 1024. We infuse the MoLT module in the last 4 layers of image and text models, meaning that $L_{s}$ is 4. $s$ is set to 4 in the FBP operations. In the loss function, we set $\alpha$ to 0.1 and $\beta$ to 0.9. To reduce the memory consumption, we train our model with mixed-precision. The Adam optimizer is employed in our method and the learning rate is set to 0.0004. Our approach is implemented with PyTorch framework and the experiments are conducted on two NVIDIA RTX 3090Ti GPUs.
Table 1: The quantitative results of our method on three benchmark datasets. w/ LM indicates whether the large models are utilized in the approach. * indicates that the original paper did not provide corresponding results, and we re-implement the method based on their open-source codes.
@@
Ours $\checkmark$ 67.0/61.9 92.44 75.10
4.3 Quantitative Results
The results of our method on the evaluated datasets are presented in Table 1. As evident from the results, we achieve state-of-the-art performance on each benchmark, demonstrating the effectiveness of our approach. Among the methods we compare with, HUSE [16] and VisualBert [6] do not utilize large models, which highlights the advantage of incorporating advanced model architectures. In contrast, other methods, such as MaPLe [7], MMBT [8], and PMF [14], are based on large models. Most large model-based methods aim to facilitate information exchange through fine-tuning strategies. Specifically, MaPLe [7] proposed a tuning module that maps linguistic embeddings to the visual branch, enhancing cross-modal interactions. Meanwhile, PMF [14] aims to disentangle vanilla prompts into different types for tuning large models, optimizing the adaptability and specificity of the prompts to various tasks. The substantial performance gap between our method and previous methods illustrates the effectiveness of our proposed technique, showcasing its potential to significantly advance the field. Our approach not only leverages the strengths of large model architectures but also introduces innovative fine-tuning mechanisms that enhance model performance across diverse datasets.
Table 2: The ablation study on SNLI-VE and UPMC-Food101 dataset. C.A. means the cross-attention in MoLT. Fusion indicates the fusion strategy in our method.
Ablation SNLI-VE UPMC-Food101
C.A. $L_{CCA}$ $M$ Fusion FBP Acc(%) Acc(%)
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 70.6 89.98
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 72.5 91.03
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 74.0 91.35
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 72.9 90.67
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 74.21 92.65
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 75.10 92.44
4.4 Ablation Study
To further investigate the effectiveness of each component in our method, we conducted an ablation study using SNLI-VE and UPMC-Food101 datasets. The results are presented in Table 2. When removing $L_{CCA}$, the performance drops dramatically, highlighting the crucial role of this training target in aligning the multimodal embeddings. This significant decrease underscores how central $L_{CCA}$ is to the model’s ability to capture and integrate cross-modal correlations. The results also demonstrate that both the cross-attention module and the learnable vector $M$ contribute positively to the final outcomes. The cross-attention module facilitates nuanced interactions between modalities, while the learnable vector $M$ adapts dynamically to enhance representation quality. When the fusion module is omitted, meaning only the robust representations are employed for predictions, there is a noticeable decrease in performance. This decline underscores the importance of the fusion module in synthesizing information from multiple modalities effectively. Correspondingly, the results also illustrate the effectiveness of the FBP module, which aims to integrate multimodal embeddings into a unified, robust representation. The FBP module’s role in enhancing the coherence and informativeness of the combined features is clearly evidenced by the performance drop observed when it is removed.
The performance gap observed between each ablative model and our final model underscores the effectiveness of our proposed method. This gap highlights how each component, from $L_{CCA}$ to the fusion and FBP modules, plays a pivotal role in achieving state-of-the-art results. Our comprehensive ablation study confirms that the synergistic integration of these components is essential for maximizing the performance of our method on the SNLI-VE dataset.
Figure 2: The visualization of some cases for our propose method and the baseline model. The examples are from the SNLI-VE dataset.
Table 3: The inference performance of our method under modality-missing and noisy scenarios. In noised-modality inference, a certain proportion of Gaussian noise is added to the original input. The baseline model indicates neither the MolT module nor robust representations were utilized for final predictions.
Models MM-IMDB UPMC-Food101 SNLI-VE
F1-micro/macro(%) Acc(%) Acc(%)
@@ 4.4 Ablation Study
Baseline(text-absence) 57.1 / 47.9 82.1 67.4
Baseline(image-absence) 59.0 / 51.2 85.2 69.3
Ours(text-absense) 62.4 / 56.7 88.9 73.2
Ours(image-absense) 63.1 / 57.0 89.2 71.0
Noised-modality Inference
Baseline(1% noise) 59.2 / 50.5 82.2 65.5
Baseline(5% noise) 56.6 / 49.7 80.9 61.0
Baseline(10% noise) 48.1 / 39.9 77.1 55.8
Baseline(20% noise) 36.9 / 27.5 52.0 39.3
Ours(1% noise) 63.0 / 56.3 87.1 68.2
Ours(5% noise) 59.9 / 54.2 86.0 62.2
Ours(10% noise) 51.1 / 43.3 79.8 56.1
Ours(20% noise) 40.0 / 31.2 53.3 39.4
@@ 4.5 Robust Inference Analysis
As our focus lies in enhancing model robustness in modality-missing inference scenarios, we conducted corresponding experiments to evaluate our approach under challenging conditions. These experiments were performed under two distinct scenarios: one involving missing modalities and the other involving noised modality signals(we manually add a certain proportion of Gaussian noise to the input). In the baseline model, neither the MolT module nor robust representations were utilized for predictions. The corresponding results are shown in Table 2.
The results indicate that each modality has a varying impact on different tasks. For instance, in movie classification, the textual modality proves to be dominant, whereas the image modality becomes more critical for tasks like visual-entailment understanding. Despite these variations, our method consistently outperforms the baseline model, demonstrating the effectiveness of our proposed approach across different tasks. Moreover, even when the model is tested with noised multimodal features, our proposed method still yields relatively strong performance, whereas the baseline model’s performance drops dramatically. This resilience to noisy inputs highlights the robustness of our approach. We further substantiate our findings through visual inspections of several cases, as shown in Figure 2. These visual inspections reveal that with the incorporation of the MolT module and robust representation learning, our model can still make accurate predictions, whereas the baseline model often fails.
@@ 5 Conclusion
In this paper, we propose a robust representation learning strategy tailored for large models. Our approach incorporates a modality latent translation module capable of translating one modality embedding to another, with the factorized bilinear pooling used to generate robust representations. Additionally, we introduce a novel fusion schema for robust representation and modality embeddings. Experiments conducted on three datasets clearly illustrate the effectiveness of our proposed method. Further analysis demonstrates the robustness of our method in modality-missing and noisy scenarios. In the future, we plan to conduct further research in robust representation learning to enhance our ability to handle modality-absence scenarios more effectively.